# Sparse Autoencoder Neural Operators: Model Recovery in Function Spaces

### v2

@@ Appendix D Appendix - Theoretical Analysis
This section provides the mathematical derivations for the training dynamics of the architectures discussed in the main text. We derive the gradient descent update rules for a single data sample, which are used to learn the dictionary parameters. These derivations consider the analytic gradient, i.e., the dictionary gradient is computed given the codes [47, 48].
See 3.1
@@
We consider the case of SAE-MLP. The gradient of the loss for $\bm{D}$} is:
@@
Hence, when ${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$, the lifted-SAE shows equivalent learning dynamics as the SAE for model recovery. ∎
See 3.2
@@
We provide the proof for the sparse generative model case when the encoder implements a proximal gradient descent type of operation in an unrolled learning setting (e.g., 6). For the SAE, the representation refinement at every encoder layer follows:
Using the assumption on the lifting-projection,
Proposition D.1 (Architectural Inference Equivalence of Lifting in SAE-CNN).
The architectural inference of a SAE-CNN is equivalent to a lifted SAE-CNN (L-SAE-CNN) if the projection is tied to the lifting operator (${\bm{P}}={\bm{L}}^{\top}$) that is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$), i.e., under the same learned model ${\bm{D}}_{c}={\bm{P}}{\bm{D}}_{L,c}$, the representation dynamics and output of the networks are equivalent.
Proof.
We provide the proof for the sparse convolutional generative model case when the encoder implements a proximal gradient descent type of operation in an unrolled learning setting (e.g., 7). For the SAE-CNN, the representation refinement for the $c$-th feature map at every encoder layer follows:
@@
Proposition D.2.
[Training Dynamics of Lifted SAE-CNN] The training dynamics of the lifted-SAE-CNN (L-SAE-CNN) ${\bm{D}}_{L,c}^{(k+1)}={\bm{D}}_{L,c}^{(k)}+\eta_{L}{\bm{L}}\left({\bm{x}}-{\bm{L}}^{\top}\left(\sum\limits_{i=1}^{C}{\bm{D}}_{L,i}^{(k)}*{\bm{z}}_{i,t}\right)\right)\star{\bm{z}}_{c,t}$, with a lifting operator $\bm{L}$} and projection $\bm{P}$}, has the effective update in the original space, expressed as: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\eta_{L}({\bm{L}}^{\top}{\bm{L}})\left({\bm{x}}-\sum\limits_{i=1}^{C}{\bm{D}}_{i}^{(k)}*{\bm{z}}_{i,t}\right)\star{\bm{z}}_{c,t}$, where ${\bm{L}}^{\top}{\bm{L}}$ acts as a preconditioner, potentially accelerating learning by inducing a more isotropic update. If the architectures satisfy the architectural inference equivalence condition (Prop.˜D.1), then the dynamics become equivalent to that of a SAE-CNN: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\eta\left({\bm{x}}-\sum\limits_{i=1}^{C}{\bm{D}}_{i}^{(k)}*{\bm{z}}_{i,t}\right)\star{\bm{z}}_{c,t}$.
@@
Proposition D.3.
[Training Dynamics of Lifted SAE-FNO] The training dynamics of the Lifted-SAE-FNO (L-SAE-FNO)’s frequency domain weights ${\bm{W}}_{L,c}^{(k+1)}={\bm{W}}_{L,c}^{(k)}+\dfrac{\eta_{L}}{\sqrt{M}}\mathcal{F}\left({\bm{L}}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\right)\odot\mathcal{F}({\bm{z}}_{c,t})$, with lifting operator $\bm{L}$} and projection $\bm{P}$}, has an effective update in the original space, expressed as: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\dfrac{\eta_{L}}{M}\left(({\bm{L}}^{\top}{\bm{L}})\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\right)\star{\bm{z}}_{c,t}$, where ${\bm{L}}^{\top}{\bm{L}}$ acts as a preconditioner, potentially accelerating learning by inducing a more isotropic update. If the architectures satisfy the architectural inference equivalence condition (Prop.˜D.6), then the dynamics are equivalent to that of an SAE-FNO: ${\bm{D}}_{c}^{\prime(k+1)}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\dfrac{1}{\sqrt{M}}\mathcal{F}^{-1}\left(\sum\limits_{i=1}^{C}\mathcal{F}({\bm{D}}_{i}^{(k)})\odot\mathcal{F}({\bm{z}}_{i,t})\right)\right)\star{\bm{z}}_{c,t}$.
@@
${\bm{W}}_{c}^{(k+1)}={\bm{W}}_{c}^{(k)}+\dfrac{\eta}{\sqrt{M}}\mathcal{F}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\odot\mathcal{F}({\bm{z}}_{c,t})$ (37)
@@
${\bm{D}}_{c}^{\prime(k+1)}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\star{\bm{z}}_{c,t}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\dfrac{1}{\sqrt{M}}\mathcal{F}^{-1}\sum\limits_{i=1}^{C}\mathcal{F}{\bm{D}}_{i}^{(k)}\odot\mathcal{F}{\bm{z}}_{i,t}\right)\star{\bm{z}}_{c,t}$ (38)
@@
Hence, when ${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$, the lifted-SAE-FNO shows equivalent learning dynamics as the standard SAE-FNO for model recovery. ∎
Proposition D.4 (Architectural Inference Equivalence of SAE-CNN and SAE-FNO).
@@
Substituting this back into eq.˜48 gives:
@@
This final expression is the SAE-FNO’s iterative encoder update for ${\bm{z}}_{c}$ in the spectral domain, where the dictionary is $\bm{W}$}. This completes the proof of architectural inference equivalence. ∎
See 3.4
@@
${\bm{W}}_{c}^{(k+1)}={\bm{W}}_{c}^{(k)}+\dfrac{\eta}{\sqrt{M}}\mathcal{F}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\odot\mathcal{F}({\bm{z}}_{c,t})$ (56)
@@
Proposition D.5 (Architectural Inference Equivalence of L-SAE-CNN and L-SAE-FNO).
@@
By construction, we incorporate the underlying model from L-SAE-FNO to L-SAE-CNN: ${\bm{W}}_{L,c}=\mathcal{F}{\bm{D}}_{L,c}$, and re-write the gradient of the encoder using Correlation and Convolution Theorems:
@@
Substituting this back into eq.˜58 gives:
@@
This final expression is the L-SAE-FNO’s iterative update for ${\bm{z}}_{c}$ in the lifted spectral domain, where the dictionary is ${\bm{W}}_{L,c}$. This completes the proof of architectural inference equivalence. ∎
Proposition D.6 (Architectural Inference Equivalence of Lifting in SAE-FNO).
The architectural inference of an SAE-FNO is equivalent to a Lifted SAE-FNO (L-SAE-FNO) if the projection is tied to the lifting operator (${\bm{P}}={\bm{L}}^{\top}$) that is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$), i.e., under the same learned model ${\bm{W}}_{c}={\bm{P}}{\bm{W}}_{L,c}$, the representation dynamics and output of the networks are equivalent.
@@
    From Proposition˜D.4, the architectural inference of an SAE-FNO is equivalent to SAE-CNN.
@@
    From Proposition˜D.1, the architectural inference of an SAE-CNN is equivalent to L-SAE-CNN.
@@
    From Proposition˜D.5, the architectural inference of a L-SAE-CNN is equivalent to L-SAE-FNO.
By the transitive property of these equivalences, we can establish a direct architectural inference equivalence between SAE-FNO and L-SAE-FNO under the same lifting-projection conditions. ∎

### v3

@@ Appendix C Appendix - Theoretical Analysis
This section provides the mathematical derivations for the training dynamics of the architectures discussed in the main text. We derive the gradient descent update rules for a single data sample, which are used to learn the dictionary parameters. These derivations consider the analytic gradient, i.e., the dictionary gradient is computed given the codes (Malézieux et al., 2022; Tolooshams and Ba, 2022).
Proposition C.1 (Architectural Inference Equivalence of Lifting in SAE).
The architectural inference of a SAE is equivalent to a lifted-SAE if the projection is tied to the lifting operator (${\bm{P}}={\bm{L}}^{\top}$), and the lifting operator is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$), i.e., under the same learned model ${\bm{D}}={\bm{P}}{\bm{D}}_{L}$, the representation dynamics and output of the networks are equivalent.
Proof.
We provide the proof for the sparse generative model case when the encoder implements a proximal gradient descent type of operation in an unrolled learning setting (e.g., 8). For the SAE, the representation refinement at every encoder layer follows:
Using the assumption on the lifting-projection,
Proposition C.2 (Training Dynamics of Lifting).
The training dynamics of the lifted-SAE (L-SAE) ${\bm{D}}_{L}^{(k+1)}={\bm{D}}_{L}^{(k)}+\eta{\bm{P}}^{T}({\bm{x}}-{\bm{P}}{\bm{D}}_{L}^{(k)}{\bm{z}}){\bm{z}}^{T}$, with lifting $\bm{L}$} and projection $\bm{P}$}, has the effective update in the original space, expressed as: ${\bm{D}}^{(k+1)}={\bm{D}}^{(k)}+\eta_{L}({\bm{L}}^{\top}{\bm{L}})({\bm{x}}-{\bm{D}}^{(k)}{\bm{z}}){\bm{z}}^{\top}$, where ${\bm{L}}^{\top}{\bm{L}}$ acts as a preconditioner, potentially accelerating learning by inducing a more isotropic update. If they satisfy the equivalent architectural inference (Proposition C.1), then the dynamics are equivalent to those of an SAE: ${\bm{D}}^{(k+1)}={\bm{D}}^{(k)}+\eta({\bm{x}}-{\bm{D}}^{(k)}{\bm{z}}){\bm{z}}^{T}$.
@@
The gradient of the loss for $\bm{D}$} is:
@@
Hence, when ${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$, the lifted-SAE shows equivalent learning dynamics as the standard SAE for model recovery. ∎
Proposition C.3 (Architectural Inference Equivalence of Lifting in SAE-CNN).
The architectural inference of a SAE-CNN is equivalent to a lifted SAE-CNN (L-SAE-CNN) if the projection is tied to the lifting operator (${\bm{P}}={\bm{L}}^{\top}$), and the lifting operator is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$), i.e., under the same learned model ${\bm{D}}_{c}={\bm{P}}{\bm{D}}_{L,c}$, the representation dynamics and output of the networks are equivalent.
@@
We provide the proof for the sparse convolutional generative model case when the encoder implements a proximal gradient descent type of operation in an unrolled learning setting (e.g., 9). For the SAE-CNN, the representation refinement for the $c$-th feature map at every encoder layer follows:
@@
Proposition C.4 (Training Dynamics of Lifted SAE-CNN).
The training dynamics of the lifted-SAE-CNN (L-SAE-CNN) ${\bm{D}}_{L,c}^{(k+1)}={\bm{D}}_{L,c}^{(k)}+\eta_{L}{\bm{L}}\left({\bm{x}}-{\bm{L}}^{\top}\left(\sum\limits_{i=1}^{C}{\bm{D}}_{L,i}^{(k)}*{\bm{z}}_{i,t}\right)\right)\star{\bm{z}}_{c,t}$, with a lifting operator $\bm{L}$} and projection $\bm{P}$}, has the effective update in the original space, expressed as: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\eta_{L}({\bm{L}}^{\top}{\bm{L}})\left({\bm{x}}-\sum\limits_{i=1}^{C}{\bm{D}}_{i}^{(k)}*{\bm{z}}_{i,t}\right)\star{\bm{z}}_{c,t}$, where ${\bm{L}}^{\top}{\bm{L}}$ acts as a preconditioner, potentially accelerating learning by inducing a more isotropic update. If the architectures satisfy the architectural inference equivalence condition (Proposition C.3), then the dynamics become equivalent to that of a SAE-CNN: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\eta\left({\bm{x}}-\sum\limits_{i=1}^{C}{\bm{D}}_{i}^{(k)}*{\bm{z}}_{i,t}\right)\star{\bm{z}}_{c,t}$.
@@
See 3.5
@@
${\bm{W}}_{c}^{(k+1)}={\bm{W}}_{c}^{(k)}+\dfrac{\eta}{\sqrt{M}}\mathcal{F}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\odot\mathcal{F}({\bm{z}}_{c,k})$ (40)
@@
${\bm{D}}_{c}^{\prime(k+1)}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\star{\bm{z}}_{c,t}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\dfrac{1}{\sqrt{M}}\mathcal{F}^{-1}\sum\limits_{i=1}^{C}\mathcal{F}{\bm{D}}_{i}^{(k)}\odot\mathcal{F}{\bm{z}}_{i,k}\right)\star{\bm{z}}_{c,t}$ (41)
@@
Hence, when ${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$, the lifted-SAE-FNO shows equivalent learning dynamics as the standard SAE-FNO for model recovery. ‘ ∎
Proposition C.5 (Architectural Inference Equivalence of SAE-CNN and SAE-FNO).
@@
Substituting this back into Equation 51 gives:
@@
This final expression is the SAE-FNO’s iterative encoder update for ${\bm{z}}_{c}$ in the spectral domain, where the dictionary is $\bm{W}$}. This completes the proof of architectural inference equivalence between SAE-CNN and SAE-FNO. ∎
Proposition C.6 (Training Dynamics of SAE-CNN and SAE-FNO).
The training dynamics of the SAE-FNO’s frequency-domain weights ${\bm{W}}_{c}^{(k+1)}={\bm{W}}_{c}^{(k)}+\dfrac{\eta_{L}}{\sqrt{M}}\mathcal{F}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\odot\mathcal{F}({\bm{z}}_{c,t})$ has an effective update in the original space, expressed as: ${\bm{D}}_{c}^{\prime(k+1)}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\star{\bm{z}}_{c,t}$. If the architectures satisfy the architectural inference equivalence condition (Proposition C.5), then the dynamics is equivalent to that of a SAE-CNN: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\eta\left({\bm{x}}-\sum\limits_{i=1}^{C}{\bm{D}}_{i}^{(k)}*{\bm{z}}_{i,t}\right)\star{\bm{z}}_{c,t}$.
@@
${\bm{W}}_{c}^{(k+1)}={\bm{W}}_{c}^{(k)}+\dfrac{\eta}{\sqrt{M}}\mathcal{F}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\odot\mathcal{F}({\bm{z}}_{c,k})$ (59)
@@
Proposition C.7 (Architectural Inference Equivalence of L-SAE-CNN and L-SAE-FNO).
@@
By construction, we incorporate the underlying model from the L-SAE-FNO to the L-SAE-CNN: ${\bm{W}}_{L,c}=\mathcal{F}{\bm{D}}_{L,c}$, and re-write the gradient of the encoder using Correlation and Convolution Theorems:
@@
Substituting this back into Equation 61 gives:
@@
This final expression is the L-SAE-FNO’s iterative update for ${\bm{z}}_{c}$ in the lifted spectral domain, where the dictionary is ${\bm{W}}_{L,c}$. This completes the proof of architectural inference equivalence between L-SAE-CNN and L-SAE-FNO. ∎
Proposition C.8 (The Architectural Inference of L-SAE-FNO reduced to SAE-FNO).
The architectural inference of an SAE-FNO is equivalent to a L-SAE-FNO if the projection is tied to the lifting operator (${\bm{P}}={\bm{L}}^{\top}$), and the lifting operator is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$), i.e., under the same learned model ${\bm{W}}_{c}={\bm{P}}{\bm{W}}_{L,c}$, the representation dynamics and output of the networks are equivalent.
@@
    From Proposition C.5, the architectural inference of an SAE-FNO is equivalent to a SAE-CNN.
@@
    From Proposition C.3, the architectural inference of a SAE-CNN is equivalent to L-SAE-CNN.
@@
    From Proposition C.7, the architectural inference of a L-SAE-CNN is equivalent to a L-SAE-FNO.
By the transitive property of these equivalences, we can establish a direct architectural inference equivalence between SAE-FNO and L-SAE-FNO under the same lifting-projection conditions. This completes the proof. ∎
Proposition C.9 (Training Dynamics of SAE-FNO).
Consider an SAE-FNO whose integral operator is implemented as a convolution, with frequency-domain weights ${\bm{W}}_{c}$ defined as the Fourier transform of the Conv-SAE’s spatial kernels, i.e., ${\bm{W}}_{c}=\mathcal{F}{\bm{D}}_{c}$. The training dynamics of the SAE-FNO are then expressed as: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\tfrac{\eta}{M}({\bm{x}}-\tfrac{1}{\sqrt{M}}\mathcal{F}^{-1}[{\textstyle\sum_{i=1}^{C}}\mathcal{F}{\bm{D}}_{i}^{(k)}\odot\mathcal{F}{\bm{z}}_{i}])\star{\bm{z}}_{c}$ where $M$ is the full number of modes, $\odot$ denotes element-wise multiplication, and $\star$ denotes correlation. If the SAE-FNO’s filters span the full frequency range and are constrained to have the same spatial support as in the Conv-SAE, then its dynamics reduce to those of the Conv-SAE (see Proposition C.6).